The Lakeside Company uses a weighted-average process costing system. The following data are available:

Beginning inventory -0-
Units started in production 24,000
Units finished during the period 17,900
Units in process at the end of the period
(complete as to materials,
¼ complete as to labor and overhead)

6,100
Cost of materials used $ 60,500
Labor and overhead costs $ 62,480

Total cost of the 17,900 units finished is:

Multiple Choice

  • $102,698.

  • $111,447.

  • $123,011.

  • $109,467.
Answer #1

Equivalent units for material = 24000 units

Equivalent units for Labor and overhead costs = 17900 + 6100*1/4 = 19425 units

Cost per unit of material = 60500 / 24000 = $2.520 per unit

cost per unit of Labor and overhead costs = 62480 / 19425 = $3.216 per unit

Total cost (2.52+3.22)*17900 = $102,698